Update October 8th, 1515: Campus Power and Network Outage(RESOLVED)

Description:

Updates for today, October 8th, are below:

Update 1515: Services have been restored for Humanities Gateway(611). At this time, our teams have resolved all issues caused by the power outage. If you have any residual issues, email oit@uci.edu.

Update 1500: Services have been restored for the Middle Earth Towers. We are waiting for final confirmation that issues affecting Humanities Gateway(611) are resolved.

Update 1400: Issues surrounding the 1st and 3rd floors of Humanities Gateway(611) are almost resolved. Field technicians are en route to assist with replacements. Engineers are in the process of confirming service restoration for the Middle Earth Towers.

Update 1200: The switch supporting Social Science Laboratory(202) has been restored, and wired services should be resolved.

Update 1100: Service has been partially restored for the 1st floor of Humanities Gateway(611). Our engineers are working to replace a switch. There are no further updates regarding Middle Earth Towers and Social Science Laboratory(202) at this time.

Update 1000: Technicians are still in the process of recovering switches in the 1st and 3rd floors of Humanities Gateway(611). A technician has been dispatched to Social Science Laboratory(202) to investigate further. Our Network Engineers are still troubleshooting wired services for the 6th and 7th floors of Middle Earth Towers, but switches supporting this area are back up.

Update 0800: Our technicians have repaired the 4th floor switch in Humanities Gateway(611) and have identified new issues for the 1st and 3rd floor. We are still investigating issues affecting wired services for the 6th and 7th floor of Middle Earth Towers. We are also working to resolve issues affecting Social Science Lab(202).

Updates for October 7th are below:

Update 1723: Service has been restored to Multipurpose Academic & Administrative Bldg (2). Our technicians are still working on Humanities Gateway (611) and the Middle Earth Towers.

Update 1715: Service has been restored to Student Services 1(105) and Multipurpose Science & Technology Bldg (415). Our technicians are currently working on replacing a switch in Humanities Gateway (611) and addressing the wired issues with the Middle Earth Towers.

Update 1630: Unfortunately, there are residual wired issues with Middle Earth Towers that aren’t fully resolved yet. However, wireless connections are online and unaffected. There is no ETA at the time.

Update 1600: Service has been restored to Middle Earth Towers and Verano Place Cyber Cafe(V34). We are still working on the remaining three affected buildings.

Update 1500: Our Network Engineers are actively investigating a confirmed issue at Middle Earth Towers 6th floor. The access point supporting the Intercollegiate Athletics Building(625) has been restored.

Update 1400: Our technicians are still working to troubleshoot and deploy affected APs. Service for the Student Center Parking Structure(SCP) has been restored.

Update 1300: The Middle Earth Towers may potentially have an affected switch from the outage. Wired connections may be unavailable at this time. Our Network Engineers are currently investigating. Access points have been configured for the Student Center Parking Structure(SCP) and the Verano Place Cyber Cafe(V34), and will be deployed soon.

Update 1100: Services have been restored to the Merage School of Business(221). Our technicians are working to install newly configured access points for MPAA(2), SSI(105), MSTB(415), and the Intercollegiate Athletics Building(625).

Update 1000: SBSG(214) has been 100% recovered.

Update 0915: Services have been restored to the 3rd and 5th floor of SBSG(214).

Update 0800: SBSG(214) switches are in the process of being recovered this morning. Techs and Engineers are 50% completed. Electrical Substation(040) and Art Culture Technology(727) have been recovered. Network Engineers will be prioritizing on access points and other residual incidents.

Updates for October 6th are below:

Update 1645: We now have an updated list of affected wireless points as well. Field technicians are working to restore service to Electrical Substation(40) and the Merage School of Business(221).

Update 1615: Our technicians have restored service for all wired connections except for the buildings listed below. Please keep in mind that some buildings not listed are still having issues with their wireless connection.

Update 1445: Our technicians are working to restore service to Law Building(4), Engineering Gateway(321), and Art Culture Technology(727) as well. Social & Behavioral Sciences Gateway(214) has been added to the list of affected buildings. Services have been restored to Education Building(3), University Extension Building(8), McGaugh Hall(503), Humanities Instructional Bldg(610), and Sprague Hall(839). Thank you for your continued patience.

Update 1300: Unfortunately, there are still residual issues with the Merage School of Business(221) and Engineering Gateway(321). Our teams are actively investigating.

Update 1230: Additionally, our technicians and engineers are also working to restore service to McGaugh Hall(503), Art Culture Technology(727), and Sprague Hall(839). Services have been restored for Electrical Substation(40), EH&S Services Facility(41), Medical Sciences C(821), and Crawford Hall(903). The list of remaining affected buildings has been updated.

Update 1100: Our technicians and engineers are working on restoring services to the Student Center(113). Services have been restored for the Merage School of Business(221), CalIT2(325), Bonney Research Lab(512), Humanities Gateway(611), and Irvine Hall(835). The list of remaining affected buildings has been updated accordingly.

Update 1000: Our engineers are also working to restore services to MSTB(2), Merage School of Business(221), and Humanities Gateway(611). Our security and network teams have replaced a failed firewall affection Bison Modular(515) as well.

Update 0900: The list of affected buildings has been updated. Technicians are currently working to restore services to Student Center A(113), McGaugh Hall(503), and the Humanities Instructional Building(610) this morning.

Updates for October 5th are below:

Update 1700: Services have been restored the following buildings: Social Sciences Laboratory(202), Engineering Gateway(321), Humanities Gateway(611), Medical Sciences B(811), and Mesa Court Otero Dorm(M35). The list of affected buildings has been updated.

Update 1630: Services are still being restored to CalIT2(325). Our technicians are still actively working to restore service to affected buildings, and we thank you for your continued patience.

Update 1415: After failed reboots, network engineers and field technicians are working with the security team to configure replacement equipment supporting Bison Modular(515). The list has been updated accordingly.

Update 1230: Service has been restored to CalIT2(325) and Beckman Laser Institute(817). The list has been updated.

Update 1220: Switches have been deployed at MSTB so wired connections should be back up. However, network engineers are still restoring service to access points, so wireless connections may still be affected at this time. Our technicians are also working on deploying a switch at Beckman Laser Institute to restore service.

Update 1040: The list of affected buildings and floors below has been updated. High priority buildings are Aldrich Hall(Floor 3), Steinhaus Hall, McGaugh Hall, MSTB, and the Beckman Laser Institute. Service to the Rockwell Engineering Center has been restored.

Update 0911: Switches in Med Sci B(811), Med Sci C(821), Beckman Laser Institute(817), and MSTB(415) are down. Service is restored to the 5th floor of Aldrich Hall.

Update 0806: Technicians are currently working on deploying switches in Aldrich Hall to restore service, ETA 1 hour. Technicians are also working on affected access points in Mesa Court, no ETA at the time.

Update 0621: Currently power outage repairs are underway but there are a number of switches and access points that are down and will need to be replaced. Here is the list of the affected buildings and affected floors:

We are receiving reports of a campus power outage. Some of the buildings affected are Rowland Hall, Anteater Learning Pavilion, Engineering Hall among others. No ETA at this time.

First reported on October 4, 2021 9:32 pm

Status: Resolved

Resolution:

Our Network Engineers, Field Technicians, and Security team have worked together to restore all affected wired and wireless devices caused by the power outage. If you have any residual issues, please email oit@uci.edu.

Contact the OIT Help Desk if you have questions or further issues.

Updated on: October 8, 2021 5:36 pm

All Service Alerts